# Cognitive Hierarchy in Day-to-day Network Flow Dynamics

### v1

@@ 4 Validation Using a Virtual Experiment
Ye et al. (2018) conducted an online route choice experiment to mimic travelers’ decision-making processes from day to day. The experiment collected 268 participants’ route choices on 26 rounds, where each round corresponded to a true calendar day. It was conducted on the well-known Braess paradox network (Figure 2) with one OD pair served by three routes: Route 1 as $1\to 3\to 5$; Route 2 as $2\to 5\to 3$; and Route 3 as $2\to 4$. The travel time on each link $a\in L$ was given by the well-known BPR function: $t_{a}(v_{a})=t_{a}^{0}\left(1+0.15\left(\frac{v_{a}}{V_{a}}\right)^{4}\right)$, with parameters marked in the figure in order as ($a$, $t_{a}^{0}$, $V_{a}$), where $t_{a}^{0}$ is the free-flow travel time and $V_{a}$ the capacity. The observed route flow (i.e., how many participants selected that route) on each day is visualized in Figure 2 with the days indexed from 0 to 25. Note the large oscillations in the trajectories in Figure 2. Readers can refer to Ye et al. (2018) for more details of the experiment.
Figure 1: Braess network used in the virtual experiment (Ye et al., 2018).
Figure 2: Collected route flow evolutions in the virtual experiment (Ye et al., 2018).
@@ 4.1 Difficulty of prevailing models without prediction in reproducing the observed pattern
A natural way of calibrating a day-to-day dynamical model is to find the optimal parameters that minimize the sum of squared error between simulated and observed flow evolution trajectories. In doing so, Ye et al. (2018) found that the prevailing day-to-day models could not produce fluctuated trajectories fitting the experimental data; see Figure 3 of that paper. (Due to the intractability of the above “simulation-based” method, they had to turn to a relaxed problem using regression analysis.) Below we give an explanation of the prevailing models’ difficulties.
Figure 3: Net flow on day $t$ for the three routes sorted by travel cost on day $t-1$ in the virtual experiment.
@@
First note that for day $t\in[1,16]$, the longest route on the day $t-1$ would by-and-large out-flow to the other two routes on the day $t$. This observation is in line with most day-to-day models’ assumption that travelers would select the routes with lower costs. However, there was no distinction between the shortest and the second-shortest routes. On quite a few days (i.e., 1, 3, 8, 9, 10, and 13), the previous day’s second-shortest route received more inflow than the shortest route. This indiscrimination between the shortest and the second-shortest routes contradicted the best-response-based day-to-day models’ underlying hypothesis that the shortest route would be the most popular (e.g., Friesz et al., 1994; He et al., 2010; Xiao et al., 2016). We argue that this indiscrimination was a consequence of strategic thinking. Specifically, a 1-step traveler might think that other travelers would go for the shortest route of the day $t-1$ and thus avoided switching to that route on the day $t$. As a result, the second-best route becomes her preference.
@@ 4.2 Calibrations of the CH-NTP dynamic
To minimize the number of parameters, we consider special cases with perfect prediction (i.e., $\hat{\gamma}=\gamma$) and $\hat{\alpha}=\alpha=1$. By doing so, we intend to demonstrate our model’s general explanation capability rather than “overfit” data using more parameters. (Relaxing these constraints will further reduce the fitting errors.) The 0-step model is exactly the so-called XYY dynamics used in Xiao et al. (2016) and Ye et al. (2018), which has only one parameter, $\gamma$, to be estimated. (The other day-to-day models (without prediction) calibrated in Ye et al. (2018) exhibited very similar (bad) performance as the 0-step model and hence are omitted here for simplicity. We aim to show how introducing cognitive hierarchy can significantly improve model interpretation.) The 1-step model has one more parameter $p^{0}$ because $p^{1}$ equals $1-p^{0}$. And the 2-step model has three parameters, namely, $\gamma$, $p^{0}$, and $p^{1}$.
@@ 4.2.1 Preparations
We conducted calibration by minimizing the root mean square error (RMSE) between the predicted flows and the ground truth:
@@
For each CH-NTP dynamic with different $|K|$, the optimal parameters that minimize the RMSE function (23) are used to calculate log-likelihoods and perform likelihood ratio tests. The log-likelihood function is derived as follows. We can first obtain a $k$-step traveler’s probability of choosing a particular route by dividing the route flow of k-step travelers by their OD demands. Denote $g^{k}\left(r_{l}^{j,(t)}\mid\gamma,p^{0},\dots,p^{|K|-1}\right)$ as the model’s predicted probability of a traveler $l$ from class $k$ choosing route $j$ on day $t$. Then the final total predicted probability, $G^{(t)}(r_{l}^{j})$, is an aggregation of all thinking steps weighted by the proportions; i.e., $G^{(t)}(r_{l}^{j})=\sum_{k}p^{k}g^{k}\left(r_{l}^{j,(t)}\mid\gamma,p^{0},\dots,p^{|K|-1}\right)$. Finally, we can form a log-likelihood function by combining all the travelers on all the routes and days:
@@
where $I\left(r_{l}^{j,(t)}\right)$ is the indicator function being 1 if traveler $l$ chooses route $j$ on day $t$ and 0 otherwise.
@@ 4.2.2 Results
The simulated flow evolutions of Route 1 with the parameters that minimized the RMSE are presented in Figures 4(a)-4(c) for $M=9$, $16$, and $25$, respectively. (The other two routes exhibit similar results and hence are omitted to conserve space.) The likelihood ratio test is reported in Table 1 with the optimal parameters.
@@
    Estimated log-likelihood using the optimal parameters that minimized RMSE function (23).
@@
First note that in all the figures, the nonhierarchical model with $|K|=1$, i.e., the conventional NTP dynamic, fails to produce a significant fluctuation pattern fitting the ground-truth trajectory; see the dotted curves. This result is in agreement with Ye et al. (2018). The corresponding flow trajectories for $M=9$, $16$, and $25$ are identical and the optimal $\gamma=0.358$.
@@
Further note that we have previously implemented a multi-class day-to-day model where each class is associated with a distinct sensitivity parameter (Zhou et al., 2020). Despite increasing the number of classes to three, this multi-class model still fails to reproduce the observed fluctuation pattern. In fact, the calibrated flow evolution from this model closely resembles the pattern seen in the $|K|=1$ case (illustrated by the red dotted curve in Figures 4(a)-4(c)). This outcome strongly suggests that our improved fitting performance is not merely a result of additional parameters, but rather stems from the enhanced explanatory capability of our proposed model.
@@
Happily, significant fluctuation patterns emerge when some travelers think one step further, as revealed by the yellow dash-dotted curves. The fitting error decreases as a consequence. Note in Figures 4(b) and 4(c) how the predicted curves faithfully capture the ground-truth trend from day 5 to 16. This is also confirmed by observing that the hierarchical model of $|K|=2$ improves the log-likelihood by 24.5 and 23.3 for $M=16$ and $25$, respectively. Moreover, the hierarchical model with $|K|=3$ almost doubles the log-likelihood improvement and exhibits better fitting performance from the very first day until day 16. All the improvements are statistically significant under the likelihood ratio test with the $p$-value less than $0.0002\%$.
@@
Fitting performance is less satisfying for days 17-25, as shown by the trailing edges of the curves in Figure 4(c). A possible reason is that as the experiment proceeded, the participants found themselves gaining little by predicting the population’s behaviors and thus exhibited somehow uniformly-random behaviors in the last eight days (see the explanations on this matter in Section 4.1). Our time-invariant parameters cannot capture such a pattern.
5 Numerical Experiments
In this section, we conduct numerical experiments to verify the previous sections’ theoretical results and highlight some features not captured by the theoretical analysis.
5.1 Evolutions towards multiple equilibria
We examine how the strategic-reasoning behavior leads the system to multiple equilibria. We use the Braess network in the virtual experiment as our first example. At the DUE, the aggregate route flow pattern is $(89.33,89.33,89.33)^{T}$. We set $p^{0}=p^{1}=0.5$ ($|K|=2$), $\hat{\alpha}=\alpha=0.3$, $\hat{\gamma}=\gamma$. Figures 5(a) and 5(b) depict the evolution trajectories in terms of aggregate route flow on each route, starting from different initial points (distinguished by different colors) for $\gamma=1.4$ and $0.2$, respectively. (Plotting two routes is sufficient as the degree of freedom is only two due to the flow conservation constraint.) At the initial point, all the routes share the same $p^{0}$ and $p^{1}$.
(a) $\gamma=1.4$
(b) $\gamma=0.2$
Figure 5: Evolutions of route flows starting from different initial route flow patterns under different $\gamma=\hat{\gamma}$ for $|K|=2$. $p^{0}=p^{1}=0.5$, $\hat{\alpha}=\alpha=0.3$.
First note in Figure 5(a) how different initial points result in different evolution trajectories and steady states under a large $\gamma$. It is interesting to see that two initial points close to each other can evolve to totally different equilibrium points, as shown by the two trajectories starting in the middle. This kind of multiple-equilibria phenomenon is rooted in travelers’ strategic-thinking behaviors rather than the asymmetry of the travel cost functions previously studied in the literature (see, e.g., Watling, 1996; Bie and Lo, 2010; Han et al., 2017). We believe this system feature would enrich behavioral explanation capability.
Things change a lot when $\gamma$ is small. With a small $\gamma$, travelers are reluctant to switch to shorter routes and tend to believe other travelers are also conservative (as $\hat{\gamma}$ is set to equal $\gamma$). As a consequence, all the initial points will gradually evolve to the DUE, as shown in Figure 5(b).
5.2 Local stability near the DUE
@@
We next test stability conditions on a larger network used in Zhang et al. (2015); see Figure 6. It contains 8 routes that connect two OD-pairs with demand $(90,90)^{T}$: Route 1 as $1\to 9\to 14$; Route 2 as $1\to 5\to 10$; Route 3 as $2\to 6\to 10$; Route 4 as $2\to 11\to 15$; Route 5 as $3\to 11\to 16$; Route 6 as $3\to 7\to 12$; Route 7 as $4\to 8\to 12$; and Route 8 as $4\to 13\to 17$. Link travel times are again given by the well-known BPR function with link numbers, free-flow travel times, and capacities marked in the figure in order as ($a$, $t_{a}^{0}$, $V_{a}$). The aggregate DUE is $(20,20,25,\allowbreak 25,25,25,20,20)^{T}$.
Figure 6: A larger testing network (from Zhang et al., 2015).
5.2.1 In the interior of the feasible route flow set
@@
We now verify Proposition 4 where predictions are inaccurate. All the parameters and initial flow patterns are the same as the above case except for the values of $\gamma$ and $\hat{\gamma}$. Figure 8 depicts whether or not the system deviates from the DUE after a small perturbation for various combinations of $\gamma$ and $\hat{\gamma}$ in a resolution of 0.05, where solid circles and crosses denote stable and unstable systems, respectively. The points on the $45^{\circ}$ solid curve represent the case of $\hat{\gamma}=\gamma$ (perfect prediction), and the solid curve lying beneath represents the case of $\hat{\gamma}=\frac{\gamma}{2}$. The approximate $\bar{\gamma}=0.79$ by Lemma 3 is also marked.
@@
First note the green shaded area where a moderately small $\hat{\gamma}$, $\frac{\gamma}{2}<\hat{\gamma}<\bar{\gamma}<\gamma$, helps stabilize the dynamic even when $\gamma$ exceeds $\bar{\gamma}$. This justifies the finding in Proposition 4 that, a mild under-prediction enlarges the stable region. In contrast, when $\hat{\gamma}$ is very small (i.e., $<\frac{\gamma}{2}$), the stable region is separated into two pieces. The eigenvalues of $\bar{Q}D_{\ast}$ that previously satisfy the stability condition when $\hat{\gamma}=\gamma<\bar{\gamma}$ fall into the unstable region due to the separation when $\hat{\gamma}<\frac{\gamma}{2}$; see the red shaded area. In addition, as expected in Remark 3, a $\hat{\gamma}>\bar{\gamma}$ will make the system permanently unstable, regardless of $\gamma$; see all the crosses above the horizontal line of $\bar{\gamma}=0.79$. Finally, for this particular numerical case, although the stable region size is reduced when over-predictions occur ($\gamma<\hat{\gamma}<\bar{\gamma}$), The eigenvalues of $\bar{Q}D_{\ast}$ still satisfy the stability condition in Proposition 4 and thus the points above the curve $\hat{\gamma}=\gamma$ are stable.
5.2.2 On the boundary of the feasible route flow set
Proposition 3 becomes invalid when the projected flow pattern is on the boundary of the feasible set. Under this circumstance, the general analytical results are invalid, and we have to resort to Fact 1 for numerically checking the stability. To test this fact, we set $|K|=2$, $\hat{\alpha}=\alpha=0.5$, $\hat{\gamma}=\gamma\in\{0.81,0.82\}$, $p^{0}=0.4,p^{1}=0.6$ and the initial route flow pattern as $\boldsymbol{x}^{k=0,(t)=0}=(0,\allowbreak 7.999,10,18,9.98,10.01,8.01,8.01)^{T}$ and $\boldsymbol{x}^{k=1,(t)=0}=(20,11.99,15,7,14.98,15.01,12.01,12.01)^{T}$, such that $JP_{0,0}$ in Fact 1 is evaluated on the boundary. From Fact 1, we have that the maximum moduli of $JP$ of $|K|=2$ for $\gamma=0.81$ and $0.82$ are 1 and 1.013, respectively. It is expected that the former is stable while the latter is unstable, as shown in Figure 9. The case of $|K|=3$ is also consistent with the numerical result of Fact 1, which is omitted here for simplicity.

### v2

@@ 4 Validation Using a Virtual Experiment
We conducted an online route choice experiment to mimic travelers’ decision-making processes from day to day (Ye, Xiao, and Yang 2018). The experiment collected 268 participants’ route choices on 26 rounds, where each round corresponded to a true calendar day. It was implemented on the well-known Braess paradox network (Figure 3) with one OD pair served by three routes: Route 1 as $1\to 3\to 5$; Route 2 as $2\to 5\to 3$; and Route 3 as $2\to 4$. The travel time on each link $a\in L$ followed the standard BPR function: $t_{a}(v_{a})=t_{a}^{0}\left(1+0.15\left(\frac{v_{a}}{V_{a}}\right)^{4}\right)$. The parameters are marked in the figure in order as ($a$, $t_{a}^{0}$, $V_{a}$), where $t_{a}^{0}$ represents the free-flow travel time and $V_{a}$ the capacity. The observed route flow (i.e., how many participants selected that route) on each day is visualized in Figure 3 with the days indexed from 0 to 25. Note the large oscillations in the trajectories in Figure 3. Readers can refer to Ye, Xiao, and Yang (2018) for more details of the experiment.
Before proceeding with the analysis, it is worth noting why we use experimental data rather than field observations. Obtaining real-world route choice data for day-to-day analysis presents significant challenges due to limitations in sensor coverage, numerous confounding factors in traffic conditions, and difficulties in extracting route-specific decisions from aggregate traffic counts. These challenges have led transportation researchers to utilize controlled experiments when studying route choice behavior (Qi et al. 2023). While experiments cannot perfectly replicate real-world conditions, our experiment described above notably narrows the gap between laboratory conditions and real-world environments in two ways: (i) it includes a large sample of 268 participants, whereas previous studies typically involve fewer than 100 participants per scenario; and (ii) it spans 26 actual calendar days, rather than condensing all rounds within a few hours as commonly practiced in laboratory settings. This allows for more natural learning and adaptation processes similar to travelers’ day-to-day experiences.
Figure 2: Braess network used in the virtual experiment (Ye, Xiao, and Yang 2018).
Figure 3: Collected route flow evolutions in the virtual experiment (Ye, Xiao, and Yang 2018).
@@ 4.1 Difficulty of prevailing models without prediction in reproducing the observed pattern
A natural way of calibrating a day-to-day dynamical model is to find the optimal parameters that minimize the RMSE between simulated and observed flow evolution trajectories. In doing so, Ye, Xiao, and Yang (2018) found that the prevailing day-to-day models could not produce fluctuated trajectories fitting the experimental data; see Figure 3 of that paper. (Due to the intractability of the above “simulation-based” method, they had to turn to a relaxed problem using regression analysis.) Below we give an explanation of the prevailing models’ difficulties.
Figure 4: Net flow on day $t$ of the routes sorted by travel cost on day $t-1$ in the virtual experiment.
@@
First note that for day $t\in[1,16]$, the longest route on the day $t-1$ would by-and-large out-flow to the other two routes on the day $t$. This observation is in line with most day-to-day models’ assumption that travelers would select the routes with lower costs. However, there was no distinction between the shortest and the second-shortest routes. On quite a few days (i.e., 1, 3, 8, 9, 10, and 13), the previous day’s second-shortest route received more inflow than the shortest route. This indiscrimination between the shortest and the second-shortest routes contradicted the best-response-based day-to-day models’ underlying hypothesis that the shortest route would be the most popular (e.g., Friesz et al. 1994, He, Guo, and Liu 2010, Xiao, Yang, and Ye 2016). We argue that this indiscrimination was a consequence of strategic thinking. Specifically, a 1-step traveler might think that other travelers would go for the shortest route of the day $t-1$ and thus avoided switching to that route on the day $t$. As a result, the second-best route becomes her preference.
@@ 4.2 Calibrations of the CH-NTP dynamic
To minimize the number of parameters, we consider special cases with perfect prediction (i.e., $\hat{\gamma}=\gamma$) and $\hat{\alpha}=\alpha=1$. By doing so, we intend to demonstrate our model’s general explanation capability rather than “overfit” data using more parameters. (Relaxing these constraints will further reduce the fitting errors.) The 0-step model is exactly the so-called XYY dynamics used in Xiao, Yang, and Ye (2016) and Ye, Xiao, and Yang (2018), which has only one parameter, $\gamma$, to be estimated. (The other day-to-day models (without prediction) calibrated in Ye, Xiao, and Yang (2018) exhibited very similar (bad) performance as the 0-step model and hence are omitted here for simplicity. We aim to show how introducing cognitive hierarchy can significantly improve model interpretation.) The 1-step model has one more parameter $p^{0}$ because $p^{1}$ equals $1-p^{0}$. And the 2-step model has three parameters, namely, $\gamma$, $p^{0}$, and $p^{1}$. Note that we limit our analysis to cognitive hierarchies with $|K|\leq 3$, as substantial experiments in behavioral game theory have found that the average thinking step across players typically lies between 1 and 2 (Camerer, Ho, and Chong 2004, Chong, Ho, and Camerer 2016).
@@ 4.2.1 Preparations
We calibrated by minimizing the RMSE between the predicted flows and the ground truth:
@@
While our day-to-day model operates at a “macroscopic” level using continuous route flow variables, it has a “microscopic” probabilistic interpretation. For each cognitive level $k$, we can divide each route flow of that level $k$ by corresponding OD demands and interpret it as the probability of choosing that route. (This can be viewed as playing mixed strategies with these probabilities in their route choice process in Krichene, Drighès, and Bayen 2015.)
With probabilistic interpretation, we can perform likelihood ratio tests. Specifically, for each CH-NTP dynamic with different $|K|$, we first obtain the optimal parameters by minimizing the RMSE in Eq. (24). With these optimized parameters, we calculate the model’s predicted probability of a traveler $l$ from class $k$ choosing route $j$ on day $t$, denoted as $g^{k}\left(r_{l}^{j,(t)}\mid\gamma,p^{0},\dots,p^{|K|-1}\right)$. Then the final total predicted probability, $G^{(t)}(r_{l}^{j})$, is an aggregation of all thinking steps weighted by the proportions; i.e., $G^{(t)}(r_{l}^{j})=\sum_{k}p^{k}g^{k}\left(r_{l}^{j,(t)}\mid\gamma,p^{0},\dots,p^{|K|-1}\right)$. Finally, we can form a log-likelihood function by combining all the travelers on all the routes and days:
@@
where $268$ denotes the number of participants in the experiment and $I\left(r_{l}^{j,(t)}\right)$ is the indicator function being 1 if traveler $l$ chooses route $j$ on day $t$ and 0 otherwise.
@@ 4.2.2 Results
The simulated flow evolutions of Route 1 with the parameters that minimized the RMSE are presented in Figures 5(a)-5(c) for $M=9$, $16$, and $25$, respectively. Each figure uses different line styles to represent the CH-NTP dynamic with varying values of $|K|\in\{1,2,3\}$, while solid lines depict the ground-truth observed data. (The other two routes exhibit similar results and hence are omitted to conserve space.) The likelihood ratio test using the optimal parameters is reported in Table 1.
@@
    Estimated log-likelihood using the optimal parameters that minimized RMSE in Eq. (24).
@@
First note that in all the figures, the nonhierarchical model with $|K|=1$, i.e., the conventional NTP dynamic, fails to produce a significant fluctuation pattern fitting the ground-truth trajectory; see the dotted curves. This result is in agreement with Ye, Xiao, and Yang (2018). The corresponding flow trajectories for $M=9$, $16$, and $25$ are identical and the optimal $\gamma=0.358$.
@@
Further note that we have implemented a multi-class day-to-day model with distinct sensitivity parameter for each class (Zhou, Xu, and Meng 2020). Despite increasing the number of classes to three, this multi-class model still fails to reproduce the observed fluctuation pattern. In fact, the calibrated flow evolution from this model closely resembles the pattern seen in the $|K|=1$ case (illustrated by the red dotted curve in Figures 5(a)-5(c)). This outcome strongly suggests that our improved fitting performance is not merely a result of additional parameters, but rather stems from the enhanced explanatory capability of our proposed model.
@@
Happily, significant fluctuation patterns emerge when some travelers think one step further, as revealed by the yellow dash-dotted curves. The fitting error decreases as a consequence. Note in Figures 5(b) and 5(c) how the predicted curves faithfully capture the ground-truth trend from day 5 to 16. This is also confirmed by observing that the hierarchical model of $|K|=2$ improves the log-likelihood by 24.5 and 23.3 for $M=16$ and $25$, respectively. Moreover, the hierarchical model with $|K|=3$ almost doubles the log-likelihood improvement and exhibits better fitting performance from the very first day until day 16. All the improvements are statistically significant under the likelihood ratio test with the $p$-value less than $0.0002\%$.
@@
Fitting performance is less satisfying for days 17-25, as shown by the trailing edges of the curves in Figure 5(c). This decline may be explained by participant behaviors in the final eight days. As the experiment continued, participants found little advantage in trying to predict how others would behave. This led them to take somehow uniformly-random behaviors. Our time-invariant parameters cannot capture such a pattern.
5 Numerical Experiments on Local Stability near the DUE
@@
In this section, we conduct numerical experiments to verify the stability conditions using a larger network in Figure 6 (Zhang et al. 2015). The network contains 8 routes that connect two OD-pairs with demand $(90,90)^{T}$: Route 1 as $1\to 9\to 14$; Route 2 as $1\to 5\to 10$; Route 3 as $2\to 6\to 10$; Route 4 as $2\to 11\to 15$; Route 5 as $3\to 11\to 16$; Route 6 as $3\to 7\to 12$; Route 7 as $4\to 8\to 12$; and Route 8 as $4\to 13\to 17$. Link travel times are again given by the well-known BPR function with link numbers, free-flow travel times, and capacities marked in the figure in order as ($a$, $t_{a}^{0}$, $V_{a}$). The aggregate DUE is $(20,20,25,\allowbreak 25,25,25,20,20)^{T}$.
Figure 6: A larger testing network (from Zhang et al. 2015).
5.1 In the interior of the feasible route flow set
@@
We next verify Proposition 4 where predictions are inaccurate. All the parameters and initial flow patterns are the same as the above case except for the values of $\gamma$ and $\hat{\gamma}$. Figure 8 shows whether the system deviates from the DUE after a small perturbation across various combinations of $\gamma$ and $\hat{\gamma}$ at a resolution of 0.05. Solid circles represent stable systems, while crosses indicate unstable systems. The points on the $45^{\circ}$ solid curve represent the case of $\hat{\gamma}=\gamma$ (perfect prediction), and the solid curve lying beneath represents the case of $\hat{\gamma}=\frac{\gamma}{2}$. The approximate $\bar{\gamma}=0.79$ by Lemma 3 is also marked.
@@
First note the green shaded area where a moderately small $\hat{\gamma}$, $\frac{\gamma}{2}<\hat{\gamma}<\bar{\gamma}<\gamma$, helps stabilize the dynamic even when $\gamma$ exceeds $\bar{\gamma}$. This justifies the finding in Proposition 4 that, a mild under-prediction enlarges the stable region. In contrast, when $\hat{\gamma}$ is very small (i.e., $<\frac{\gamma}{2}$), the stable region is separated into two pieces. The eigenvalues of $\bar{Q}D_{\ast}$ that previously satisfy the stability condition when $\hat{\gamma}=\gamma<\bar{\gamma}$ fall into the unstable region due to the separation when $\hat{\gamma}<\frac{\gamma}{2}$; see the red shaded area. In addition, as expected in Remark 4, a $\hat{\gamma}>\bar{\gamma}$ will make the system permanently unstable, regardless of $\gamma$; see all the crosses above the horizontal line of $\bar{\gamma}=0.79$. Finally, for this particular numerical example, the stable region diminishes when over-predictions occur ($\gamma<\hat{\gamma}<\bar{\gamma}$). Nevertheless, the eigenvalues of $\bar{Q}D_{\ast}$ still satisfy the stability condition in Proposition 4, which explains why the points above the curve $\hat{\gamma}=\gamma$ remain stable.
5.2 On the boundary of the feasible route flow set
Proposition 3 becomes invalid when the projected flow pattern is on the boundary of the feasible set. Under this circumstance, the general analytical results are invalid, and we have to resort to Fact 1 for numerically checking the stability. To test this fact, we set $|K|=2$, $\hat{\alpha}=\alpha=0.5$, $\hat{\gamma}=\gamma\in\{0.81,0.82\}$, $p^{0}=0.4,p^{1}=0.6$ and the initial route flow pattern as $\boldsymbol{x}^{k=0,(t)=0}=(0,\allowbreak 7.999,10,18,9.98,10.01,8.01,8.01)^{T}$ and $\boldsymbol{x}^{k=1,(t)=0}=(20,11.99,15,7,14.98,15.01,12.01,12.01)^{T}$, such that $JP_{0,0}$ in Fact 1 is evaluated on the boundary. From Fact 1, we have that the maximum moduli of $JP$ of $|K|=2$ for $\gamma=0.81$ and $0.82$ are 1 and 1.013, respectively. This indicates that the former should exhibit stability while the latter should demonstrate instability. Figure 9 confirms these expectations, with different colors distinguishing the two $\gamma$ values and distinct line styles marking the route flows of different classes. The case of $|K|=3$ is also consistent with the numerical result of Fact 1, which is omitted here for simplicity.